### v1

# Accessing proton number fluctuations in limited regions of phase space

Abstract
In heavy ion collisions particle distributions fluctuate from event to event. It is interesting to study local fluctuations of a specific particle specie, e.g. baryons, in the transverse plane. Fluctuations of the harmonic flow provide an integrated measure of such fluctuations. An alternative approach is to study proton number fluctuations measured in a relatively narrow azimuthal angle window. Due to the transverse flow, particles registered in a narrow azimuthal angle window are emitted predominantly from a region of the fireball located at the same azimuthal angle. Similarly, it is informative to measure the covariance of the number of protons emitted in two opposite windows of azimuthal angles. The main difficulty lies in distinguishing effects of local baryon density fluctuations in the transverse plane from global particle density and collective flow fluctuations. These global volume and flow fluctuations can be estimated using some reference particles, e.g. charged mesons. Reliable estimates of such fluctuations can be constructed both for the second factorial cumulant of the proton number in an azimuthal angle window and for the covariance between two such windows. The simultaneous measurement of both quantities provides a sensitive probe of local baryon number fluctuations superimposed on top of a fluctuating fireball background density.
ultrarelativistic nuclear collisions, baryon number fluctuations, event-by-event fluctuations
I Introduction
Heavy ion collisions at relativistic energies are studied to investigate the properties and the dynamics of strongly interacting matter at extreme conditions [1, 2, 3, 4, 5]. The most important goal is to explore experimentally the phase diagram of Quantum Chromodynamics (QCD) at finite temperature and baryon density. In particular, a major part of the experimental program in relativistic heavy-ion collisions [6, 7, 8, 9, 10, 11, 12] is devoted to the search for a possible critical point in the QCD phase diagram at finite temperature and density [13, 14, 15, 16, 17, 18, 19]. The dynamics of a heavy ion collision is typically modeled as a complex process involving an initial formation of a dense fireball, its collective expansions, and final hadron production [20, 21, 22, 23, 24].
In this context, specific measures of event by event fluctuations in particle production are of particular importance [25, 26]. Proton number fluctuations are expected to increase significantly in the vicinity of the QCD critical point [27, 28]. Measuring moments of the proton multiplicity distribution as a function of the collision energy could reveal the approach of the critical point in the course of the dynamics in a heavy-ion collision at a specific beam energy. Typically one considers the (net)-proton multiplicity distribution in a given rapidity range. Near the critical point one expects critical slowing down of the build up of correlations [29, 30, 31], which may limit the correlations length to scales smaller than the transverse size of the fireball. As a result, measurements that include all particles emitted tend to average out any critical fluctuations.
Conversely, by considering particles observed within a narrow window of azimuthal angles one restricts the emission region to a localized part the fireball. This is due to the presence of a strong transverse flow in the fireball. Particles flying a given azimuthal angle are emitted preferentially from a region of the fireball located at the same azimuthal angle. Studying baryon fluctuations in the transverse plane is also relevant to understand baryon stopping and entropy production in the initial state [32, 33]. For example, one may explore the differences between the initial entropy and baryon densities. A toy model based on this idea is used in this paper to demonstrate the sensitivity of the proposed methods to such fluctuations.
Measuring the proton multiplicity distribution in a restricted azimuthal angle window is challenging due to reduced statistics but also because the emission probability fluctuates strongly event by event due to fluctuations of the collective flow. For a fixed azimuthal angle acceptance window, the number of registered protons varies with the harmonic flow in each event. Moreover, global proton number fluctuations can arise from overall rapidity density fluctuations.
In this paper, we propose a method to estimate the contribution from such global emission probability fluctuations to proton multiplicity distributions using fluctuations of some reference particles, such as charged mesons. With an accurate account for volume fluctuations contributions, one can identify a possible excess in proton fluctuations indicating the presence of local baryon density fluctuations.
We also study the covariance of the multiplicity of protons emitted in two opposite azimuthal angle windows. In this case, the volume fluctuation contribution is found to account for most of the measured covariance. Local fluctuations in the baryon distribution can thus be identified by jointly measuring the proton second factorial cumulant and the covariance, and comparing the results to volume fluctuation estimates.
In Sec. II, I describe a simple model based on relativistic hydrodynamics and the factorization breaking coefficient between the harmonic flow of protons and mesons is calculated. In Sec III, the second factorial moment of the proton distribution is studied along with estimates of the corresponding contribution from volume fluctuations. The covariance of the proton numbers emitted in opposite directions is discussed in Sec. IV. It is shown that volume fluctuations can largely explain most of the observed covariance.
II Model of proton fluctuations
The goal of this study is to explore the feasibility of measuring specific fluctuations of a chosen particle of interest. These particles of interest may constitute a particular subset of all particles emitted in a collision, where one expects to observe some specific collective behavior or correlations distinct from that of the reference particles. Such particles may include those carrying a conserved charge, e.g. strange particles, baryons, or particles containing heavy quarks. Harmonic flow coefficients of these particles of interest are routinely investigated both experimentally and theoretically. Observables that involve higher-order azimuthal asymmetries for particles of interest and reference particles have also been considered [34].
In this work, we focus on fluctuations of baryons relative to reference particles. Due to experimental considerations these are protons. Specifically, we study fluctuations of protons in comparison to charged mesons (pions and kaons). Although this case serves as a concrete example, the methods developed are broadly applicable. For instance, the approach could readily be extended to consider net-proton fluctuations. The aim is to identify fluctuations in the distribution of particles of interest that occur on top of those associated with general event by event fluctuations of the bulk. These background fluctuations are always present due to variations in multiplicity, rapidity distributions, transverse flow, and harmonic flow. The main question is how to identify additional fluctuations specific to the particles of interest (e.g. protons, or net protons) on top of the baseline provided by the bulk properties of the fireball.
The collision dynamics is modeled using the 2+1dimensional relativistic viscous hydrodynamic code MUSIC [35, 36, 37] with shear viscosity $\eta/s=0.08$. An equation of state for quark-gluon plasma at finite baryon density is used [38]. The initial distribution of entropy in the transverse plane is generated via a Monte-Carlo Glauber model [39] :
$s(x,y)\propto\sum_{i}(1-\alpha+\alpha N^{coll}_{i})g(x-x_{i},y-y_{i})\ ,$ (1)
where the sum runs over all the participant nucleons. $(x_{i},y_{i})$ denotes the position of the participant nucleon $i$ in the transverse plane, $N^{coll}_{i}$ is the number of collisions involving that participant nucleon, $g=\frac{1}{2\pi\sigma_{w}^{2}}exp\left(-\frac{(x-x_{i})^{2}+(y-y_{i})^{2}}{2\sigma_{w}}\right)$. The parameters used are $\alpha=0.18$, $\sigma_{w}=0.5$fm. The initial baryon density at central rapidity incorporates an additional factor
$\rho_{B}(x,y)\propto\sum_{i}\Gamma_{i}(1-\alpha+\alpha N^{coll}_{i})g(x-x_{i},y-y_{i})\ ,$ (2)
where $\Gamma_{i}$ is a random variable sampled from a gamma distribution with mean one and width $\sigma_{B}$. Note that the choice of the initial model is not unique; alternative models such as TRENTO [40] or models using different schemes for the fluctuations of the initial entropy and baryon densities could also be employed [32, 33]. The chosen model serves to illustrate the sensitivity of final observables to differences in the initial entropy and baryon densities in the fireball. For each parameter set, 600 initial condition are generated and evolved with the hydrodynamic code. Final state particles are generated using the ISS code [35, 36, 37], with an oversampling by a factor $20000$. The simulations are performed for Au+Au collisions at $\sqrt{s}=19.6$MeV in the $0-5$% centrality class. The normalization factors for the initial entropy (Eq. 1) and baryon (Eq. 2) densities are adjusted to reproduce the average charged particle and proton densities at midrapidity.
Figure 1: Factorization breaking coefficient (7) between the harmonic flow of protons and mesons. The solid, dashed and dotted lines represent the results for the harmonic flow of order 2,3, and 4, respectively.
In each event, the number of particles $N$ emitted within the selected kinematic interval fluctuates. A measure of the multiplicity distribution can be obtained calculating the factorial moments
$F_{n}=\langle N(N-1)\dots(N-n+1)\rangle\ ,$ (3)
[…]
$C_{1}$ $=$ $F_{1}$
$C_{2}$ $=$ $F_{2}-F_{1}^{2}$
$C_{3}$ $=$ $F_{3}-3F_{2}F_{1}+2F_{1}^{2}$
$C_{4}$ $=$ $F_{4}-4F_{3}F_{1}+12F_{2}F_{1}^{2}-3F_{2}^{2}-6F_{1}^{4}\ .$ (4)
In the presence of collective flow, the density of particles emitted at central rapidity ($|y|<1$) in a given event can be written as
$\frac{d{\cal N}}{d\phi}=\frac{\cal N}{2\pi}\left(1+2\sum_{n=1}^{\infty}V_{n}(p)e^{in\phi}\right)\ \ ,$ (5)
where $\cal N$ is the rapidity density of emitted particles and $V_{n}=v_{n}e^{in\Phi_{n}}$ is the complex harmonic flow vector of order $n$. $v_{n}=|V_{n}|$ is the flow magnitude and $\Phi_{n}$ is the flow angle orientation. The average multiplicity is $C_{1}=F_{1}=\langle{\cal N}\rangle=\langle{N}\rangle$. The harmonic flow coefficients are
$v_{n}\{2\}^{2}=\langle\frac{1}{N(N-1)}\sum_{i\neq j}e^{i\ n(\phi_{i}-\phi_{j})}\rangle\ =\langle V_{n}V_{n}^{\star}\rangle.$ (6)
In general, the proton and meson emission densities are correlated, though subtle differences may arise due to a different mass-dependent flow response. Moreover, additional fluctuations in the initial proton distribution may cause the initial eccentricities of protons to deviate from those of the bulk, enhancing decorrelation between the proton and meson harmonic flows. Fig. 1 shows the factorization breaking coefficient between mesons and protons
$r_{n}^{meson,proton}=\frac{\langle V_{n}^{meson}V_{n}^{meson\ \star}\rangle}{\sqrt{\langle V_{n}^{meson}V_{n}^{meson\ \star}\rangle\langle V_{n}^{proton}V_{n}^{proton\ \star}\rangle}}\ .$ (7)
As the strength of the baryon fluctuations increases, the deviation of the factorization breaking coefficients from one becomes more pronounced. The interpretation of the factorization breaking coefficient requires a comparison to realistic model calculations. In general, however, a rapid change in the value of the factorization breaking coefficient as a function of collision energy could signal an increase in baryon fluctuations related to the approach of the QCD critical point.
III Factorial cumulants of the proton distribution
If the correlation length of baryon fluctuations is much smaller than the transverse size of the fireball, then proton number fluctuations integrated over all angles consist of a sum of contributions from several, partially independent domains. To increase the sensitivity to local fluctuations of the baryon number, without averaging over the whole transverse plane, it is advantageous to study proton number fluctuations within a restricted azimuthal angle window. In the presence of the transverse collective flow, particles emitted within a restricted window of azimuthal angles predominately originate from a region of the fireball located on the same side of the fireball as the observation window (Fig. 2). Therefore, measuring the proton multiplicity distribution in a narrow azimuthal angle window provides sensitivity to the baryon number fluctuation in the associated emission region only. Local baryon number fluctuations can thus be observed. Conversely, fluctuations of the baryon number in regions located at opposite sides of the fireball are not correlated and would give no contribution to the covariance of the proton number observed at two opposite azimuthal angles.
Figure 2: View of the fireball in the transverse plane. The transverse flow (red arrows) correlates the position of the baryon excess in a region of the transverse plane with the number of protons emitted in corresponding azimuthal angle interval. The number of protons emitted back to back is mostly correlated by the global collective flow. Local fluctuations of the baryon number do not significantly correlate the number of protons emitted in opposite directions.
Fluctuations in the number of particles emitted in a restricted azimuthal angle window have a contribution from collective effects. Such fluctuations can be understood as volume fluctuations due to event by event fluctuations of the size of the effective fireball region emitting particles in the selected observation window. Estimating this contribution is essential to find a signature of additional, local fluctuations in the distribution of particles of interest.
Volume fluctuations have two contributions. The first one is an overall event by event fluctuation of the density of emitted particles. In Eq. (5) it is captured as fluctuations of the particle density $\cal N$ [42] and fluctuations of the strength of the harmonic flow coefficients $v_{n}$. The second contribution comes from the random orientation of the chosen azimuthal angle window with respect to the flow vector angles $\Phi_{n}$.
In general, it is difficult to provide an explicit formula for the contribution of volume and flow fluctuations, without imposing specific assumptions on the underlying distributions. In this paper, a practical formula is proposed to estimate these contributions. For this purpose the observed distribution of reference particles is used, with the assumption that fluctuations in the number of reference particles are dominated by volume and flow fluctuations. In the following, the term volume fluctuations is used to describe all such fluctuations present in the distribution of reference particles, e.g. volume fluctuations, flow fluctuations, transverse momentum conservation. By measuring these fluctuations for reference particles, one can obtain an empirical estimate of volume fluctuations and use it to infer its magnitude for the particles of interest.
Figure 3: Second factorial cumulant of the proton number as a function of the width of the azimuthal angle window. The case where the baryon distribution follows the entropy density is shown with the solid line, the results with additional fluctuations of width $\sigma_{B}=0.5$ in the baryon density are represented with the dotted line. The dashed and dash-dotted lines show the corresponding estimated contribution of volume fluctuations to the proton number cumulant (Eq. 16).
Let us write the joint distribution for the multiplicities of particles of interest (protons) and reference particles (mesons) in the chosen rapidity and azimuthal angle intervals as
$P(N,M)=\frac{{\cal N}^{N}e^{-{\cal N}}}{N!}\frac{{\cal M}^{M}e^{-{\cal M}}}{M!}G({\cal N},{\cal M})\ .$ (8)
When the particle emission is modeled as an independent emission from a given fluctuating volume, the random variables $\cal N$ and $\cal M$ can be interpreted as the effective emission volumes. Although the exact form of the distribution distribution $G({\cal N},{\cal M})$ is not known, it is reasonable to assume a strong correlation between the emission volumes $\cal N$ and $\cal M$. A minimal model is to assume a two-dimensional Gaussian distribution
$G({\cal N},{\cal M})=\frac{1}{2\pi\sqrt{det(\Sigma)}}exp\left(-\frac{1}{2}D^{T}\Sigma^{-1}D)\right)$ (9)
where
$D=\begin{pmatrix}\cal N-\langle{\cal N}\rangle\\ \cal M-\langle{\cal M}\rangle\end{pmatrix}$ (10)
and
$\Sigma=\begin{pmatrix}\sigma_{\cal N}^{2}&Cov_{\cal N,\cal M}\\ Cov_{\cal N,\cal M}&\sigma_{\cal M}^{2}\end{pmatrix}$ (11)
is the covariance matrix for the variables $\cal N$ and $\cal M$. The parameters of the covariance matrix $\Sigma$ can be estimated from the measured proton and meson distributions.
$\sigma_{\cal N}^{2}$ $=$ $\langle N(N-1)\rangle-\langle N\rangle^{2}=C_{2}^{proton}$
$\sigma_{\cal M}^{2}$ $=$ $\langle M(M-1)\rangle-\langle M\rangle^{2}=C_{2}^{meson}$
$Cov_{\cal N,\cal M}$ $=$ $\langle NM\rangle-\langle N\rangle\langle M\rangle=Cov(p,m)\ ,$ (12)
where $N$ and $M$ are the multiplicities of protons and mesons observed in a given azimuthal angle window of width $\Delta\phi$.
Figure 4: Second factorial cumulant of the proton number, in an azimuthal angle window of width $\Delta\phi=\pi/4$, as a function of the strength of the fluctuations in baryon deposition (solid line). The dashed line shows the estimated contribution of volume fluctuations to the proton factorial cumulant (Eq. 16).
The variance of the proton distribution at fixed meson volume is
[…]
The above formula has been proposed to calculate partial correlations and variances of observables under constraints on other observables [43]. It is equivalent to the assumption of a linear relations between the meson and the proton emission volumes [44]. The proton emission volume, corrected for fluctuations of the meson volume, is given by
${\cal N}|_{corrected}={\cal N}-\frac{Cov_{\cal N,\cal M}}{\sigma_{\cal M}^{2}}\left({\cal M}-\langle{\cal M}\rangle\right)\ .$ (14)
The second proton factorial cumulant under the condition of a constant meson emission volume is
$C_{2}^{proton}|_{\cal M}=C_{2}^{proton}-\frac{Cov(p,m)^{2}}{C_{2}^{meson}}$ (15)
In the following we use the quantity
$C_{2}^{vol}=\frac{Cov(p,m)^{2}}{C_{2}^{meson}}$ (16)
as an estimate of the contribution to the proton second factorial cumulant originating from fluctuations common to both particle species—such as volume and flow fluctuations. Any deviation of the measured second factorial cumulant for protons from the baseline prediction (16) signals the presence of additional, species-specific fluctuations. In particular, this difference could indicate local fluctuations of the baryon number not shared with the meson distribution. Such specific fluctuations could originate from differences in the initial densities or in the dynamics.
The formula for the volume fluctuation contribution can be applied quite generally, for other choices of the particles of interest an reference particles. The formula is not limited to applications for particles measured in azimuthal angle intervals. It works for any choice of the kinematic acceptance as well, e.g. for all azimuthal angles. Please note, that the condition of fixed meson emission volume is not equivalent to the condition of fixed meson multiplicity. The latter condition leads to a correction of the form $\frac{Cov(p,m)^{2}}{\langle M^{2}\rangle-\langle M\rangle^{2}}$ [43]. Please also note, that an assumption of perfect proportionality of the emission volumes for protons and mesons would lead a volume contributing for the factorial cumulant of order $n$ of the form $\left(\frac{\langle N\rangle}{\langle M\rangle}\right)^{n}C_{n}^{meson}$
Figure 5: Ratio of the second factorial cumulant of the proton number, in an azimuthal angle window of width $\Delta\phi=\pi/4$, to the estimate of the contribution from volume fluctuations as a function of the strength of the fluctuations of the number of stopped baryons . The solid line is for protons with $0<p_{T}<3GeV$ and the dashed line for protons with $0.8GeV<p_{T}<3GeV$.
Fig. 3 shows the second factorial cumulant of the proton multiplicity as a function of the azimuthal angle window width $\Delta\phi$. For the calculation (solid line in Fig. 3) using similar shapes of initial entropy (Eq. 1) and baryon (Eq. 1 with $\sigma_{B}=0$) densities the resulting second factorial cumulant follows closely the contribution from volume fluctuations (dashed line). When the initial baryon density in the transverse plane has additional fluctuations (Eq. 1 with $\sigma_{B}=0.5$), the proton factorial cumulant is much larger (dotted line). The corresponding contribution from volume fluctuations (dash-dotted line) cannot explain the magnitude of $C_{2}$. Unless additional assumptions on the nature of these decorrelations can be justified, the formula (16) represents an optimal experimental estimate for the contribution from volume fluctuations. The relative excess of the proton factorial cumulant over its volume contribution decreases with the width of the azimuthal angle window, which is an effect of averaging out baryon fluctuations over larger parts of the fireball.
In Fig. 4 the second factorial cumulant of the proton number is plotted as a function of the strength of fluctuations in the baryon number deposition $\sigma_{B}$, for a narrow azimuthal angle window of $\Delta\phi=\pi/4$. The solid line shows the total cumulant, while the dashed line represents the estimated volume fluctuation contribution according to Eq. (16). A clear enhancement of the proton cumulant is observed as the strength of baryon fluctuations increases, exceeding the expectations based solely on volume fluctuations. This analysis demonstrates that measuring factorial cumulants in narrow angular intervals allows for the identification of localized fluctuations in the distribution of baryons, beyond any global collective fluctuations captured captured by meson distributions.
It means that an experimental signal of possible interesting fluctuations in proton distributions could be identified at least qualitatively on top of standard volume fluctuations. The response of the emitted particles to the collective flow depends on their mass and transverse momentum. In Fig. 5 is shown the ratio of the second factorial cumulant to the corresponding volume fluctuations estimate. Two case are compared with proton transverse momenta $30<p_{T}<$ 3GeV (solid line) and $0.8$GeV$<p_{t}<3$GeV (dashed line). Protons with a larger transverse momenta have relatively more fluctuations compared to the reference particles. To obtain the most reliable estimate of the volume fluctuation contribution it is preferable to use a possibly broad kinematic acceptance in the meson and proton transverse momenta. In that way, the correlation between the proton and meson emission volumes is maximal.
For higher cumulants the contribution from volume fluctuations can be estimated using the relation between the effective proton and meson emission volumes (Eq. 14). The contribution from volume fluctuations is
$C_{n}^{vol}=\left(\frac{Cov(p,m)}{C_{2}^{meson}}\right)^{n}{C_{n}^{meson}}\ .$ (17)
In the model used in this paper both $C_{3}$ and $C_{4}$ as well as the corresponding volume contributions $C_{3}^{vol}$ and $C_{4}^{vol}$ are consistent with zero or very small. Of course, calculating higher factorial cumulants and volume contribution remains an important measurement in cases when critical fluctuations are expected [28], and will be considered in future studies, using a more elaborate model of baryon fluctuations.
IV Back to back covariance for protons
The emission of particles in opposite directions can be correlated due to the collective flow and conservation of the total transverses momentum. On the other hand, local fluctuations in the baryon density in the fireball are not expected to significantly correlate the emission back to back (Fig. 2). The covariance of the proton number in two back to back azimuthal angle intervals is expected to be dominated by the contribution from volume fluctuations. Such volume fluctuations can be estimated using measurements for reference particles - charged mesons.
Figure 6: The covariance of the number of protons in two back to back azimuthal angle windows as a function of the window width. The results without additional fluctuations in the baryon density are represented with the solid line and the results with additional fluctuations are represented with the dotted line. The dashed and dash-dotted lines represent the estimates of the contribution of volume fluctuations to the covariance (Eq. 25).
Figure 7: The covariance of the number of protons in two back to back azimuthal angle windows as a function of the strength of fluctuations of the number of stopped baryons (solid line). The dashed line shows the estimate of the contribution of volume fluctuations to the covariance (Eq. 25).
To estimate the volume fluctuation contribution for the covariance of protons emitted back to back let us consider the joint probability distribution to observe a number of protons $N_{0}$, $N_{\pi}$ and mesons $M_{0}$, $M_{\pi}$ in two azimuthal angle intervals centered at $\phi=0$ and $\pi$
$P_{4}(N_{0},N_{\pi},M_{0},M_{\pi})=\frac{{\cal N}_{0}^{N_{0}}e^{-{\cal N}_{0}}}{N_{0}!}\frac{{\cal N}_{\pi}^{N_{\pi}}e^{-{\cal N}_{\pi}}}{N_{\pi}!}$
$\frac{{\cal M}_{0}^{M_{0}}e^{-{\cal M}_{0}}}{M_{0}!}\frac{{\cal M}_{\pi}^{M_{\pi}}e^{-{\cal M}_{\pi}}}{M_{\pi}!}G_{4}({\cal N}_{0},{\cal M}_{0},{\cal N}_{\pi},{\cal M}_{\pi})\ ,$ (18)
where
$G_{4}({\cal N}_{0},{\cal M}_{0},{\cal N}_{\pi},{\cal M}_{\pi})=$
$\frac{1}{(2\pi)^{2}\sqrt{det(\Sigma_{4})}}exp\left(-\frac{1}{2}D_{4}^{T}\Sigma_{4}^{-1}D_{4})\right)$ (19)
is a 4-dimensional Gaussian distribution for the emission volumes of protons and mesons in the two azimuthal angle intervals. Here,
$D_{4}=\begin{pmatrix}{\cal N}_{0}-\langle{\cal N}_{0}\rangle\\ {\cal M}_{0}-\langle{\cal M}_{0}\rangle\\ {\cal N}_{\pi}-\langle{\cal N}_{\pi}\rangle\\ {\cal M}_{\pi}-\langle{\cal M}_{\pi}\rangle\end{pmatrix}$ (20)
and the covariance matrix between the emission volumes is
$\Sigma_{4}=\begin{pmatrix}\sigma_{\cal N}^{2}&Cov_{{\cal N},{\cal M}}&Cov_{{\cal N}_{0},{\cal N}_{\pi}}&Cov_{{\cal N}_{0},{\cal M}_{\pi}}\\ Cov_{{\cal N},{\cal M}}&\sigma_{\cal M}^{2}&Cov_{{\cal N}_{0},{\cal M}_{\pi}}&Cov_{{\cal M}_{0},{\cal M}_{\pi}}\\ Cov_{{\cal N}_{0},{\cal N}_{\pi}}&Cov_{{\cal N}_{0},{\cal M}_{\pi}}&\sigma_{\cal M}^{2}&Cov_{{\cal N},{\cal M}}\\ Cov_{{\cal N}_{0},{\cal M}_{\pi}}&Cov_{{\cal M}_{0},{\cal M}_{\pi}}&Cov_{{\cal N},{\cal M}}&\sigma_{\cal N}^{2}\end{pmatrix}\ .$ (21)
The upper left and lower right $2\times 2$ submatrices match the covariance matrix for the two-dimensional Gaussian distribution in Eq. (9). There are three additional covariances between emission volumes in opposite azimuthal angle intervals
$Cov_{{\cal N}_{0},{\cal N}_{\pi}}$ $=$ $\langle N_{0}N_{\pi}\rangle-\langle N\rangle^{2}=Cov_{bb}^{proton}$
$Cov_{{\cal N}_{0},{\cal M}_{\pi}}$ $=$ $\langle N_{0}M_{\pi}\rangle-\langle N\rangle\langle M\rangle=Cov_{bb}(p,m)$
$Cov_{{\cal M}_{0},{\cal M}_{\pi}}$ $=$ $\langle M_{0}M_{\pi}\rangle-\langle M\rangle^{2}=Cov_{bb}^{meson}\ .$ (22)
The covariance of the proton distribution at fixed meson emission volumes ${\cal M}_{0}$ and ${\cal M}_{\pi}$ is
$Cov_{bb}^{proton}|_{{\cal M}_{0},{\cal M}_{\pi}}=Cov_{bb}^{proton}-Cov_{bb}^{vol}$ (23)
and the variance of the proton distribution is
$C_{2}^{proton}|_{{\cal M}_{0},{\cal M}_{\pi}}=C_{2}^{proton}-C_{2}^{vol,4D}\ ,$ (24)
with
$Cov_{bb}^{vol}=\frac{-2Cov(p,m)Cov_{bb}(p,m)C_{2}^{meson}-Cov_{bb}^{meson}\left(Cov(p,m)^{2}+Cov_{bb}(p,m)^{2}\right)}{(C_{2}^{meson}-Cov_{bb}^{meson})(C_{2}^{meson}+Cov_{bb}^{meson})}$ (25)
and
$C_{2}^{vol,4D}=\frac{-2Cov_{bb}^{meson}Cov_{bb}(p,m)Cov_{bb}(p,m)+C_{2}^{meson}\left(Cov(p,m)^{2}+Cov_{bb}(p,m)^{2}\right)}{(C_{2}^{meson}-Cov_{bb}^{meson})(C_{2}^{meson}+Cov_{bb}^{meson})}\ .$ (26)
I have verified that the expression (26) for the volume fluctuations contribution to the proton factorial cumulant yields results numerically very close to the simpler expression in Eq. (16). On the other hand, the expression (25) will be used in the following as an estimate of the volume fluctuation contribution to the covariance of proton multiplicities observed in two back to back azimuthal angle windows.
Figure 8: The dynamical correlation coefficient of the proton multiplicities in two back to back azimuthal angle intervals as a function of the strength of fluctuations in the number of stopped baryons (solid line). The dashed line represents the estimate of the contribution of volume fluctuations (Eq. 25). The dotted line represents corresponding correlation coefficient for meson.
The covariance of proton multiplicities observed at opposite azimuthal angles is nearly independent on the strength of local fluctuations in the deposition of baryons, for $\Delta\phi\leq\pi$ (compare solid and dash-dotted lines in Fig. 6). The proton covariance closely matches the contribution from volume fluctuations (dashed and dotted lines in Fig. 6). In narrow azimuthal angle windows ($\Delta\phi=\pi/4$) the covariance of proton multiplicities is almost entirely explained by volume fluctuations; even for the most extreme baryon deposition fluctuations assumed (Fig. 7).
As the strength of baryon density fluctuations increases, the second factorial cumulant of the proton number increases beyond the volume fluctuation contribution (Fig. 4), while the covariance of proton multiplicities in two azimuthal angle intervals remains approximately constant (Fig. 7). Consequently, the dynamical correlation coefficient (defined with the second factorial cumulant in the denominator)
$\rho_{bb}=\frac{\langle N_{0}N_{\pi}\rangle}{\langle N(N-1)\rangle-\langle N\rangle^{2}}$ (27)
decreases with increasing baryon density fluctuations (Fig. 8). The corresponding meson correlation coefficient remains approximately constant.
It is preferable to measure separately both the covariance and the second factorial cumulant for protons, along with the associated volume fluctuation contributions and not just the dynamical correlation coefficient (Eq. 27). An excess of the proton second factorial cumulant over the volume fluctuation baseline, combined with a proton covariance consistent with volume fluctuations, provides strong evidence for short-range fluctuations in the baryon distribution in the transverse plane.
V Summary and Outlook
The paper presents a method to search for local fluctuations in the fireball produced in relativistic heavy-ion collisions. In particular, I study the fluctuations of a specific type of particles of interest (protons) with respect to reference particles (charged mesons). The method allows to partly separate global fluctuations of the particle emission from fluctuations on smaller scales. In order to be sensitive to local fluctuations in the transverse plane, I propose to consider fluctuations of particle emission in restricted windows of azimuthal angle. Due to the presence of a strong transverse flow in heavy-ion collisions, particles emitted at a given azimuthal angle originate predominantly from a region of the fireball located at that same angle. Strong global fluctuations in the particle emission occur due to fluctuations of the collective flow event by event, e.g. fluctuations of the magnitude and orientation of the harmonic flow vectors. Such global volume fluctuation contributions can be estimated calculating fluctuations of the meson distribution and scaling accordingly.
Complementary information is obtained from the covariance of the number of protons emitted in two opposite intervals of azimuthal angle. The dominant contribution to the covariance comes from global volume fluctuations. It can be estimated from measurable covariances involving mesons and protons. The proton covariance is weakly dependent on the strength of local fluctuations in the baryon deposition. Thus, a novel method is proposed, measuring simultaneously the second factorial cumulant for protons and the covariance and comparing it to the corresponding volume fluctuation baseline. Such a procedure can identify the presence of short range fluctuations in the baryon densities that may carry information about the underlying microscopic dynamics.
In particular, the method discussed here could be used in the search for fluctuations in baryon density in the vicinity of the QCD critical point [28, 19]. In order to validate the method based on the study of proton distributions in restricted intervals of azimuthal angle further studies are planned. These involve : the calculation for net baryons and net protons, using a model with nontrivial higher order proton cumulants, including effects of global baryon conservation, including effects of baryon diffusion, and making predictions for different collisions energies using realistic, 3+1-dimensional hydrodynamic simulations. Finally, the dependence of the baryon factorial cumulants on the width of the azimuthal angle window could be potentially sensitive to the value of the correlations length of the underlying baryon fluctuations.
Acknowledgments
The author thanks Adam Bzdak for useful discussions. This research was supported by the AGH University of Science and Technology and by the Polish National Science Centre grant: 2023/51/B/ST2/01625.

### v2

@@
 In heavy ion collisions particle distributions fluctuate from event to event.
-It is interesting to study local fluctuations of a specific particle specie, e.g.
-baryons, in the transverse plane.
+It is interesting to study local fluctuations of a specific particle species, e.g., baryons, in the transverse plane.
 Fluctuations of the harmonic flow provide an integrated measure of such fluctuations.
 An alternative approach is to study proton number fluctuations measured in a relatively narrow azimuthal angle window.
-Due to the transverse flow, particles registered in a narrow azimuthal angle window are emitted predominantly from a region of the fireball located at the same azimuthal angle.
+Due to the transverse flow, particles registered in a narrow azimuthal angle window are emitted predominantly from a region located at the same side of the fireball as the observation window.
 Similarly, it is informative to measure the covariance of the number of protons emitted in two opposite windows of azimuthal angles.
 The main difficulty lies in distinguishing effects of local baryon density fluctuations in the transverse plane from global particle density and collective flow fluctuations.
-These global volume and flow fluctuations can be estimated using some reference particles, e.g.
-charged mesons.
+These global density and flow fluctuations can be estimated using some reference particles, e.g., charged mesons.
 Reliable estimates of such fluctuations can be constructed both for the second factorial cumulant of the proton number in an azimuthal angle window and for the covariance between two such windows.
@@
 Measuring moments of the proton multiplicity distribution as a function of the collision energy could reveal the approach of the critical point in the course of the dynamics in a heavy-ion collision at a specific beam energy.
-Typically one considers the (net)-proton multiplicity distribution in a given rapidity range.
-Near the critical point one expects critical slowing down of the build up of correlations [29, 30, 31], which may limit the correlations length to scales smaller than the transverse size of the fireball.
-As a result, measurements that include all particles emitted tend to average out any critical fluctuations.
+Typically one considers the proton multiplicity distribution in a given rapidity interval.
+The baryon correlation range in rapidity has been studied [29, 30] using the dependence of the scaled cumulants on the width of the rapidity window.
+Existing data show no dependence up to the rapidity window width $|\Delta y|<0.5$.
+It would be interesting to estimate the range of baryon correlations in the transverse direction as well.
+Near the critical point one expects critical slowing down of the build up of correlations [31, 32, 33], which may cause the correlation lengths to never reach the whole transverse size of the fireball.
+As a result, measurements that include all emitted particles tend to average out any possible critical fluctuations.
 Conversely, by considering particles observed within a narrow window of azimuthal angles one restricts the emission region to a localized part the fireball.
+Such a measurement could provide information on baryon fluctuations in a limited region relevant to the actual scale of baryon correlations in the system [34].
 This is due to the presence of a strong transverse flow in the fireball.
-Particles flying a given azimuthal angle are emitted preferentially from a region of the fireball located at the same azimuthal angle.
+Particles flying at a given azimuthal angle are emitted preferentially from a region of the fireball located around that angle.
 Studying baryon fluctuations in the transverse plane is also relevant to understand baryon stopping and entropy production in the initial state [35, 36].
-For example, one may explore the differences between the initial entropy and baryon densities.
+For example, one may explore differences between the initial entropy and baryon densities.
 A toy model based on this idea is used in this paper to demonstrate the sensitivity of the proposed methods to such fluctuations.
-Measuring the proton multiplicity distribution in a restricted azimuthal angle window is challenging due to reduced statistics but also because the emission probability fluctuates strongly event by event due to fluctuations of the collective flow.
+Measuring the proton multiplicity distribution in a restricted azimuthal angle window is challenging due to reduced statistics.
+Note that experimental data on proton cumulants have been obtained in small rapidity windows [29], which suggests that measurements in smaller azimuthal angle intervals are possible as well.
+Another problem is that the emission probability fluctuates strongly event by event due to fluctuations of the collective flow.
 For a fixed azimuthal angle acceptance window, the number of registered protons varies with the harmonic flow in each event.
 Moreover, global proton number fluctuations can arise from overall rapidity density fluctuations.
-In this paper, we propose a method to estimate the contribution from such global emission probability fluctuations to proton multiplicity distributions using fluctuations of some reference particles, such as charged mesons.
-With an accurate account for volume fluctuations contributions, one can identify a possible excess in proton fluctuations indicating the presence of local baryon density fluctuations.
-We also study the covariance of the multiplicity of protons emitted in two opposite azimuthal angle windows.
+In this paper, I propose a method to estimate the contribution from such global emission probability fluctuations to proton multiplicity distributions using fluctuations of some reference particles, such as charged mesons.
+With an accurate account for volume fluctuations contributions, one can identify a possible excess in proton fluctuations, indicating the presence of local baryon density fluctuations.
+I also study the covariance of the multiplicity of protons emitted in two opposite azimuthal angle windows.
 In this case, the volume fluctuation contribution is found to account for most of the measured covariance.
@@
 II, I describe a simple model based on relativistic hydrodynamics and the factorization breaking coefficient between the harmonic flow of protons and mesons is calculated.
-In Sec III, the second factorial moment of the proton distribution is studied along with estimates of the corresponding contribution from volume fluctuations.
+In Sec III, the second factorial moment of the proton number distribution is studied along with estimates of the corresponding contribution from volume fluctuations.
 The covariance of the proton numbers emitted in opposite directions is discussed in Sec.
@@
 II Model of proton fluctuations
-The goal of this study is to explore the feasibility of measuring specific fluctuations of a chosen particle of interest.
+The goal of this study is to explore the feasibility of the measurement of specific fluctuations of a chosen particle of interest.
 These particles of interest may constitute a particular subset of all particles emitted in a collision, where one expects to observe some specific collective behavior or correlations distinct from that of the reference particles.
-Such particles may include those carrying a conserved charge, e.g.
-strange particles, baryons, or particles containing heavy quarks.
+Such particles may include those carrying a conserved charge, e.g., strange particles, baryons, or particles containing heavy quarks.
 Harmonic flow coefficients of these particles of interest are routinely investigated both experimentally and theoretically.
 Observables that involve higher-order azimuthal asymmetries for particles of interest and reference particles have also been considered [37].
-In this work, we focus on fluctuations of baryons relative to reference particles.
+In this work, I focus on fluctuations of baryons relative to reference particles.
 Due to experimental considerations these are protons.
-Specifically, we study fluctuations of protons in comparison to charged mesons (pions and kaons).
+Specifically, I study fluctuations of protons in comparison to charged mesons (pions and kaons).
 Although this case serves as a concrete example, the methods developed are broadly applicable.
-For instance, the approach could readily be extended to consider net-proton fluctuations.
-The aim is to identify fluctuations in the distribution of particles of interest that occur on top of those associated with general event by event fluctuations of the bulk.
+The aim is to identify fluctuations in the distribution of particles of interest that occur on top of those associated with general event by event fluctuations of the bulk of the matter.
 These background fluctuations are always present due to variations in multiplicity, rapidity distributions, transverse flow, and harmonic flow.
-The main question is how to identify additional fluctuations specific to the particles of interest (e.g.
-protons, or net protons) on top of the baseline provided by the bulk properties of the fireball.
-The collision dynamics is modeled using the 2+1dimensional relativistic viscous hydrodynamic code MUSIC [35, 36, 37] with shear viscosity $\eta/s=0.08$.
+The main question is how to identify additional fluctuations specific to the particles of interest (in this case protons) on top of the baseline provided by the bulk properties of the fireball.
+The collision dynamics is modeled using the 2+1-dimensional relativistic viscous hydrodynamic code MUSIC [38, 39, 40] with shear viscosity $\eta/s=0.08$.
 An equation of state for quark-gluon plasma at finite baryon density is used [41].
@@
 Final state particles are generated using the ISS code [38, 39, 40], with an oversampling by a factor $20000$.
-The simulations are performed for Au+Au collisions at $\sqrt{s}=19.6$MeV in the $0-5$% centrality class.
+This procedure effectively washes out nonflow correlations in the model simulations.
+The simulations are performed for Au+Au collisions at $\sqrt{s_{NN}}=19.6$GeV in the 050-5% centrality class.
 The normalization factors for the initial entropy (Eq.
@@
 2) densities are adjusted to reproduce the average charged particle and proton densities at midrapidity.
-Figure 1: Factorization breaking coefficient (7) between the harmonic flow of protons and mesons.
+Figure 1: Factorization breaking coefficient (7) between the harmonic flow of protons and mesons as a function of the strength of fluctuations in baryon deposition.
 The solid, dashed and dotted lines represent the results for the harmonic flow of order 2,3, and 4, respectively.
@@
 $F_{n}=\langle N(N-1)\dots(N-n+1)\rangle\ ,$ (3)
-[…]
+where $\langle\dots\rangle$ is the average over the events. The corresponding lowest order factorial cumulants are [44]
 $C_{1}$ $=$ $F_{1}$
@@
 The harmonic flow coefficients are
-$v_{n}\{2\}^{2}=\langle\frac{1}{N(N-1)}\sum_{i\neq j}e^{i\ n(\phi_{i}-\phi_{j})}\rangle\ =\langle V_{n}V_{n}^{\star}\rangle.$ (6)
+$v_{n}\{2\}^{2}=\left\langle\frac{1}{N(N-1)}\sum_{i\neq j}e^{i\ n(\phi_{i}-\phi_{j})}\right\rangle\ =\langle V_{n}V_{n}^{\star}\rangle.$ (6)
 In general, the proton and meson emission densities are correlated, though subtle differences may arise due to a different mass-dependent flow response.
-Moreover, additional fluctuations in the initial proton distribution may cause the initial eccentricities of protons to deviate from those of the bulk, enhancing decorrelation between the proton and meson harmonic flows.
+Moreover, additional fluctuations in the initial baryon density profile may cause the initial eccentricities of protons to deviate from those of the rest of the matter, enhancing the decorrelation between the proton and meson harmonic flows.
 Fig.
-1 shows the factorization breaking coefficient between mesons and protons
-$r_{n}^{meson,proton}=\frac{\langle V_{n}^{meson}V_{n}^{meson\ \star}\rangle}{\sqrt{\langle V_{n}^{meson}V_{n}^{meson\ \star}\rangle\langle V_{n}^{proton}V_{n}^{proton\ \star}\rangle}}\ .$ (7)
+1 shows the factorization breaking coefficient between mesons and protons,
+$r_{n}^{meson,proton}=\frac{\langle V_{n}^{meson}V_{n}^{proton\ \star}\rangle}{\sqrt{\langle V_{n}^{meson}V_{n}^{meson\ \star}\rangle\langle V_{n}^{proton}V_{n}^{proton\ \star}\rangle}}\ .$ (7)
 As the strength of the baryon fluctuations increases, the deviation of the factorization breaking coefficients from one becomes more pronounced.
@@
 If the correlation length of baryon fluctuations is much smaller than the transverse size of the fireball, then proton number fluctuations integrated over all angles consist of a sum of contributions from several, partially independent domains.
-To increase the sensitivity to local fluctuations of the baryon number, without averaging over the whole transverse plane, it is advantageous to study proton number fluctuations within a restricted azimuthal angle window.
+To increase the sensitivity to local fluctuations of the baryon number, without averaging over the whole transverse plane, it is advantageous to study proton number fluctuations within a restricted azimuthal angle window in the laboratory frame [34].
 In the presence of the transverse collective flow, particles emitted within a restricted window of azimuthal angles predominately originate from a region of the fireball located on the same side of the fireball as the observation window (Fig.
 2).
-Therefore, measuring the proton multiplicity distribution in a narrow azimuthal angle window provides sensitivity to the baryon number fluctuation in the associated emission region only.
+A similar effect has been discussed in the context of femtoscopic correlations between pairs of particles emitted with small relative momenta.
+The effective size of the emission region is much smaller than the size of the whole fireball [45].
+The experimentally observed femtoscopic radii decrease with increasing average transverse momentum of the pair [46].
+This means that the effective emission region decreases when narrowing the azimuthal angle interval of the pair.
+The effect is also visible in dynamical model simulations tracing the emission region for particles emitted in a narrow azimuthal angle interval [47].
+Such a kinematic effect of the transverse flow means that measuring the proton multiplicity distribution in a narrow azimuthal angle window provides sensitivity to the baryon number fluctuation in the associated emission region only.
 Local baryon number fluctuations can thus be observed.
@@
 Figure 2: View of the fireball in the transverse plane.
-The transverse flow (red arrows) correlates the position of the baryon excess in a region of the transverse plane with the number of protons emitted in corresponding azimuthal angle interval.
-The number of protons emitted back to back is mostly correlated by the global collective flow.
-Local fluctuations of the baryon number do not significantly correlate the number of protons emitted in opposite directions.
-Fluctuations in the number of particles emitted in a restricted azimuthal angle window have a contribution from collective effects.
-Such fluctuations can be understood as volume fluctuations due to event by event fluctuations of the size of the effective fireball region emitting particles in the selected observation window.
-Estimating this contribution is essential to find a signature of additional, local fluctuations in the distribution of particles of interest.
-Volume fluctuations have two contributions.
-The first one is an overall event by event fluctuation of the density of emitted particles.
-In Eq.
-(5) it is captured as fluctuations of the particle density $\cal N$ [42] and fluctuations of the strength of the harmonic flow coefficients $v_{n}$.
-The second contribution comes from the random orientation of the chosen azimuthal angle window with respect to the flow vector angles $\Phi_{n}$.
-In general, it is difficult to provide an explicit formula for the contribution of volume and flow fluctuations, without imposing specific assumptions on the underlying distributions.
+The trans- verse flow (red arrows) induces a correlation between the position of the baryon excess in a region of the transverse plane and the number of protons emitted in the corresponding azimuthal angle interval.
+The multiplicities of protons emitted back to back are mostly correlated due to the presence of the global collective flow.
+On the other hand, additional local fluctuations of the baryon number do not induce any significant correlation between the multiplicities of protons emitted in opposite directions.
+In this paper, volume fluctuations are understood as event by event fluctuations of the total expected number of protons registered in a given acceptance window.
+In the case of measurements at all azimuthal angles, such volume fluctuations are determined by the number of stopped baryons in a given event [48].
+On the other hand, fluctuations in the number of particles emitted in a restricted azimuthal angle window have a contribution from collective flow effects.
+For massive particles emitted in a restricted transverse momentum interval, event by event fluctuations of the transverse collective flow induce fluctuations in the expected number of particles in the acceptance window.
+The random orientation of the azimuthal flow angles contributes to multiplicity fluctuations of particles observed in an azimuthal interval (fixed in the laboratory frame).
+In the following, the term volume fluctuations is used to describe all such fluctuations present in the distribution of reference particles, e.g., density fluctuations, azimuthal flow fluctuations, transverse flow fluctuations.
+All such fluctuations will be understood as volume fluctuations in this paper, as they are due to event by event fluctuations of the size of the effective region of the fireball emitting particles in the selected acceptance window.
+Estimating all such background contributions (overall density fluctuations, transverse flow fluctuations, azimuthal flow strength and orientation fluctuations) is essential in order to find a signature of any additional, local fluctuations in the distribution of particles of interest.
+In the specific case of the fluctuations of total baryon num- ber, explicit formulas for such background fluctuations can be written [48, 30].
+In general, it is difficult to provide an explicit formula for the contribution of volume fluctuations for the moments of the multiplicity of particles emitted in a restricted acceptance window, without imposing specific assumptions on the underlying distributions.
+For multiplicity distributions in a restricted azimuthal angle interval the dominant background contribution from the elliptic flow could be explicitly subtracted [34].
+However, any interesting fluctuations that one seeks would also contribute to the elliptic flow.
 In this paper, a practical formula is proposed to estimate these contributions.
-For this purpose the observed distribution of reference particles is used, with the assumption that fluctuations in the number of reference particles are dominated by volume and flow fluctuations.
-In the following, the term volume fluctuations is used to describe all such fluctuations present in the distribution of reference particles, e.g.
-volume fluctuations, flow fluctuations, transverse momentum conservation.
+For this purpose the observed distribution of reference particles is used, with the assumption that fluctuations in the number of reference particles are dominated by volume fluctuations.
 By measuring these fluctuations for reference particles, one can obtain an empirical estimate of volume fluctuations and use it to infer its magnitude for the particles of interest.
@@
 16).
-Let us write the joint distribution for the multiplicities of particles of interest (protons) and reference particles (mesons) in the chosen rapidity and azimuthal angle intervals as
+One can write the joint distribution for the multiplicitiy $N$ of particles of interest (protons) and the multiplicity $M$ of reference particles (mesons) $M$ in the chosen rapidity and azimuthal angle intervals as
 $P(N,M)=\frac{{\cal N}^{N}e^{-{\cal N}}}{N!}\frac{{\cal M}^{M}e^{-{\cal M}}}{M!}G({\cal N},{\cal M})\ .$ (8)
-When the particle emission is modeled as an independent emission from a given fluctuating volume, the random variables $\cal N$ and $\cal M$ can be interpreted as the effective emission volumes.
-Although the exact form of the distribution distribution $G({\cal N},{\cal M})$ is not known, it is reasonable to assume a strong correlation between the emission volumes $\cal N$ and $\cal M$.
-A minimal model is to assume a two-dimensional Gaussian distribution
+[…]
 $G({\cal N},{\cal M})=\frac{1}{2\pi\sqrt{det(\Sigma)}}exp\left(-\frac{1}{2}D^{T}\Sigma^{-1}D)\right)$ (9)
 where
-$D=\begin{pmatrix}\cal N-\langle{\cal N}\rangle\\ \cal M-\langle{\cal M}\rangle\end{pmatrix}$ (10)
+[…]
 and
 $\Sigma=\begin{pmatrix}\sigma_{\cal N}^{2}&Cov_{\cal N,\cal M}\\ Cov_{\cal N,\cal M}&\sigma_{\cal M}^{2}\end{pmatrix}$ (11)
-is the covariance matrix for the variables $\cal N$ and $\cal M$. The parameters of the covariance matrix $\Sigma$ can be estimated from the measured proton and meson distributions.
-$\sigma_{\cal N}^{2}$ $=$ $\langle N(N-1)\rangle-\langle N\rangle^{2}=C_{2}^{proton}$
+[…]
+$\sigma_{\cal N}^{2}$ $=$ $\langle N(N-1)\rangle-\langle N\rangle^{2}=C_{2}$
 $\sigma_{\cal M}^{2}$ $=$ $\langle M(M-1)\rangle-\langle M\rangle^{2}=C_{2}^{meson}$
@@
 where $N$ and $M$ are the multiplicities of protons and mesons observed in a given azimuthal angle window of width $\Delta\phi$.
-Figure 4: Second factorial cumulant of the proton number, in an azimuthal angle window of width $\Delta\phi=\pi/4$, as a function of the strength of the fluctuations in baryon deposition (solid line).
+Figure 4: Second factorial cumulant of the proton number, in an azimuthal angle window of width $\Delta\phi=\pi/4$, as a function of the strength of fluctuations in baryon deposition (solid line).
 The dashed line shows the estimated contribution of volume fluctuations to the proton factorial cumulant (Eq.
@@
 The variance of the proton distribution at fixed meson volume is
-[…]
+$\sigma_{\cal N}|_{\cal M}^{\ \ \ 2}=\sigma_{\cal N}^{2}-\frac{Cov_{\cal N,\cal M}^{2}}{\sigma_{\cal M}^{2}}\ .$ (13)
 The above formula has been proposed to calculate partial correlations and variances of observables under constraints on other observables [49].
@@
 The second proton factorial cumulant under the condition of a constant meson emission volume is
-$C_{2}^{proton}|_{\cal M}=C_{2}^{proton}-\frac{Cov(p,m)^{2}}{C_{2}^{meson}}$ (15)
-In the following we use the quantity
-$C_{2}^{vol}=\frac{Cov(p,m)^{2}}{C_{2}^{meson}}$ (16)
-as an estimate of the contribution to the proton second factorial cumulant originating from fluctuations common to both particle species—such as volume and flow fluctuations.
+$C_{2}|_{\cal M}=C_{2}-\frac{Cov(p,m)^{2}}{C_{2}^{meson}}$ (15)
+In the following I use the quantity
+$C_{2}^{vol}=\frac{Cov(p,m)^{2}}{C_{2}^{meson}}\$(16)
+as an estimate of the contribution to the proton second factorial cumulant originating from fluctuations common to both particle species, such as volume fluctuations.
 Any deviation of the measured second factorial cumulant for protons from the baseline prediction (16) signals the presence of additional, species-specific fluctuations.
@@
 Such specific fluctuations could originate from differences in the initial densities or in the dynamics.
-The formula for the volume fluctuation contribution can be applied quite generally, for other choices of the particles of interest an reference particles.
+The formula for the volume fluctuation contribution can be applied quite generally, for other choices of the particles of interest and reference particles.
 The formula is not limited to applications for particles measured in azimuthal angle intervals.
-It works for any choice of the kinematic acceptance as well, e.g.
-for all azimuthal angles.
-Please note, that the condition of fixed meson emission volume is not equivalent to the condition of fixed meson multiplicity.
-The latter condition leads to a correction of the form $\frac{Cov(p,m)^{2}}{\langle M^{2}\rangle-\langle M\rangle^{2}}$ [43].
-Please also note, that an assumption of perfect proportionality of the emission volumes for protons and mesons would lead a volume contributing for the factorial cumulant of order $n$ of the form $\left(\frac{\langle N\rangle}{\langle M\rangle}\right)^{n}C_{n}^{meson}$
-Figure 5: Ratio of the second factorial cumulant of the proton number, in an azimuthal angle window of width $\Delta\phi=\pi/4$, to the estimate of the contribution from volume fluctuations as a function of the strength of the fluctuations of the number of stopped baryons .
+It works for any choice of the kinematic acceptance as well, e.g., for a fixed rapidity interval at all azimuthal angles.
+Please note that the condition of fixed meson emission volume is not equivalent to the condition of fixed meson multiplicity.
+The latter condition leads to a correction of the form [49]
+$\frac{Cov(p,m)^{2}}{\langle M^{2}\rangle-\langle M\rangle^{2}}\ .$ (17)
+Please also note, that an assumption of perfect proportionality of the emission volumes for protons and mesons would lead a volume contributing for the factorial cumulant of order $n$ of the form
+$\left(\frac{\langle N\rangle}{\langle M\rangle}\right)^{n}C_{n}^{meson}\ .$ (18)
+Figure 5: Ratio of the second factorial cumulant of the proton number, in an azimuthal angle window of width $\Delta\phi=\pi/4$, to the estimate of the contribution from volume fluctuations as a function of the strength of fluctuations of the number of stopped baryons .
 The solid line is for protons with $0<p_{T}<3GeV$ and the dashed line for protons with $0.8GeV<p_{T}<3GeV$.
@@
 The corresponding contribution from volume fluctuations (dash-dotted line) cannot explain the magnitude of $C_{2}$.
-Unless additional assumptions on the nature of these decorrelations can be justified, the formula (16) represents an optimal experimental estimate for the contribution from volume fluctuations.
 The relative excess of the proton factorial cumulant over its volume contribution decreases with the width of the azimuthal angle window, which is an effect of averaging out baryon fluctuations over larger parts of the fireball.
+However, with decreasing width of the azimuthal angle interval the statistical uncertainty increases.
 In Fig.
@@
 It means that an experimental signal of possible interesting fluctuations in proton distributions could be identified at least qualitatively on top of standard volume fluctuations.
-The response of the emitted particles to the collective flow depends on their mass and transverse momentum.
-In Fig.
-5 is shown the ratio of the second factorial cumulant to the corresponding volume fluctuations estimate.
-Two case are compared with proton transverse momenta $30<p_{T}<$ 3GeV (solid line) and $0.8$GeV$<p_{t}<3$GeV (dashed line).
-Protons with a larger transverse momenta have relatively more fluctuations compared to the reference particles.
-To obtain the most reliable estimate of the volume fluctuation contribution it is preferable to use a possibly broad kinematic acceptance in the meson and proton transverse momenta.
-In that way, the correlation between the proton and meson emission volumes is maximal.
+[…]
 For higher cumulants the contribution from volume fluctuations can be estimated using the relation between the effective proton and meson emission volumes (Eq.
@@
 $C_{n}^{vol}=\left(\frac{Cov(p,m)}{C_{2}^{meson}}\right)^{n}{C_{n}^{meson}}\ .$ (19)
-In the model used in this paper both $C_{3}$ and $C_{4}$ as well as the corresponding volume contributions $C_{3}^{vol}$ and $C_{4}^{vol}$ are consistent with zero or very small.
+In the model used in this paper both $C_{3}$ and $C_{4}$ as well as the corresponding volume contributions $C_{3}^{vol}$ and $C_{4}^{vol}$ are consistent with zero or are very small.
 Of course, calculating higher factorial cumulants and volume contribution remains an important measurement in cases when critical fluctuations are expected [28], and will be considered in future studies, using a more elaborate model of baryon fluctuations.
 IV Back to back covariance for protons
-The emission of particles in opposite directions can be correlated due to the collective flow and conservation of the total transverses momentum.
-On the other hand, local fluctuations in the baryon density in the fireball are not expected to significantly correlate the emission back to back (Fig.
+The emission of particles in opposite directions can be correlated due to the collective flow and conservation of the total transverse momentum.
+On the other hand, local fluctua- tions of the baryon density in the fireball are not expected to induce significant correlations in the multiplicities of protons emitted back to back (Fig.
 2).
@@
 The results without additional fluctuations in the baryon density are represented with the solid line and the results with additional fluctuations are represented with the dotted line.
-The dashed and dash-dotted lines represent the estimates of the contribution of volume fluctuations to the covariance (Eq.
-25).
+The dashed and dash-dotted lines represent the respective estimates of the contribution of volume fluctuations to the covariance (Eq.
+27).
 Figure 7: The covariance of the number of protons in two back to back azimuthal angle windows as a function of the strength of fluctuations of the number of stopped baryons (solid line).
 The dashed line shows the estimate of the contribution of volume fluctuations to the covariance (Eq.
-25).
-To estimate the volume fluctuation contribution for the covariance of protons emitted back to back let us consider the joint probability distribution to observe a number of protons $N_{0}$, $N_{\pi}$ and mesons $M_{0}$, $M_{\pi}$ in two azimuthal angle intervals centered at $\phi=0$ and $\pi$
+27).
+To estimate the volume fluctuation contribution for the covariance of protons emitted back to back one can consider the joint probability distribution to observe a number of protons $N_{0}$, $N_{\pi}$ and mesons $M_{0}$, $M_{\pi}$ in two azimuthal angle intervals centered at $\phi=0$ and $\pi$
 $P_{4}(N_{0},N_{\pi},M_{0},M_{\pi})=\frac{{\cal N}_{0}^{N_{0}}e^{-{\cal N}_{0}}}{N_{0}!}\frac{{\cal N}_{\pi}^{N_{\pi}}e^{-{\cal N}_{\pi}}}{N_{\pi}!}$
@@
 $\frac{1}{(2\pi)^{2}\sqrt{det(\Sigma_{4})}}exp\left(-\frac{1}{2}D_{4}^{T}\Sigma_{4}^{-1}D_{4})\right)$ (21)
-is a 4-dimensional Gaussian distribution for the emission volumes of protons and mesons in the two azimuthal angle intervals. Here,
+is a four-dimensional Gaussian distribution for the emission volumes of protons and mesons in the two azimuthal angle intervals. Here,
 $D_{4}=\begin{pmatrix}{\cal N}_{0}-\langle{\cal N}_{0}\rangle\\ {\cal M}_{0}-\langle{\cal M}_{0}\rangle\\ {\cal N}_{\pi}-\langle{\cal N}_{\pi}\rangle\\ {\cal M}_{\pi}-\langle{\cal M}_{\pi}\rangle\end{pmatrix}$ (22)
 and the covariance matrix between the emission volumes is
-$\Sigma_{4}=\begin{pmatrix}\sigma_{\cal N}^{2}&Cov_{{\cal N},{\cal M}}&Cov_{{\cal N}_{0},{\cal N}_{\pi}}&Cov_{{\cal N}_{0},{\cal M}_{\pi}}\\ Cov_{{\cal N},{\cal M}}&\sigma_{\cal M}^{2}&Cov_{{\cal N}_{0},{\cal M}_{\pi}}&Cov_{{\cal M}_{0},{\cal M}_{\pi}}\\ Cov_{{\cal N}_{0},{\cal N}_{\pi}}&Cov_{{\cal N}_{0},{\cal M}_{\pi}}&\sigma_{\cal M}^{2}&Cov_{{\cal N},{\cal M}}\\ Cov_{{\cal N}_{0},{\cal M}_{\pi}}&Cov_{{\cal M}_{0},{\cal M}_{\pi}}&Cov_{{\cal N},{\cal M}}&\sigma_{\cal N}^{2}\end{pmatrix}\ .$ (21)
+$\Sigma_{4}=\begin{pmatrix}\sigma_{\cal N}^{2}&Cov_{{\cal N},{\cal M}}&Cov_{{\cal N}_{0},{\cal N}_{\pi}}&Cov_{{\cal N}_{0},{\cal M}_{\pi}}\\ Cov_{{\cal N},{\cal M}}&\sigma_{\cal M}^{2}&Cov_{{\cal N}_{0},{\cal M}_{\pi}}&Cov_{{\cal M}_{0},{\cal M}_{\pi}}\\ Cov_{{\cal N}_{0},{\cal N}_{\pi}}&Cov_{{\cal N}_{0},{\cal M}_{\pi}}&\sigma_{\cal M}^{2}&Cov_{{\cal N},{\cal M}}\\ Cov_{{\cal N}_{0},{\cal M}_{\pi}}&Cov_{{\cal M}_{0},{\cal M}_{\pi}}&Cov_{{\cal N},{\cal M}}&\sigma_{\cal N}^{2}\end{pmatrix}\$.
+(23)
 The upper left and lower right $2\times 2$ submatrices match the covariance matrix for the two-dimensional Gaussian distribution in Eq.
 (9).
-There are three additional covariances between emission volumes in opposite azimuthal angle intervals
-$Cov_{{\cal N}_{0},{\cal N}_{\pi}}$ $=$ $\langle N_{0}N_{\pi}\rangle-\langle N\rangle^{2}=Cov_{bb}^{proton}$
+There are three additional covariances between emission volumes in opposite azimuthal angle intervals:
+$Cov_{{\cal N}_{0},{\cal N}_{\pi}}$ $=$ $\langle N_{0}N_{\pi}\rangle-\langle N\rangle^{2}=Cov_{bb}$
 $Cov_{{\cal N}_{0},{\cal M}_{\pi}}$ $=$ $\langle N_{0}M_{\pi}\rangle-\langle N\rangle\langle M\rangle=Cov_{bb}(p,m)$
@@
 The covariance of the proton distribution at fixed meson emission volumes ${\cal M}_{0}$ and ${\cal M}_{\pi}$ is
-$Cov_{bb}^{proton}|_{{\cal M}_{0},{\cal M}_{\pi}}=Cov_{bb}^{proton}-Cov_{bb}^{vol}$ (23)
+$Cov_{bb}|_{{\cal M}_{0},{\cal M}_{\pi}}=Cov_{bb}-Cov_{bb}^{vol}$ (25)
 and the variance of the proton distribution is
-$C_{2}^{proton}|_{{\cal M}_{0},{\cal M}_{\pi}}=C_{2}^{proton}-C_{2}^{vol,4D}\ ,$ (24)
+$C_{2}|_{{\cal M}_{0},{\cal M}_{\pi}}=C_{2}-C_{2}^{vol,4D}\ ,$ (26)
 with
-$Cov_{bb}^{vol}=\frac{-2Cov(p,m)Cov_{bb}(p,m)C_{2}^{meson}-Cov_{bb}^{meson}\left(Cov(p,m)^{2}+Cov_{bb}(p,m)^{2}\right)}{(C_{2}^{meson}-Cov_{bb}^{meson})(C_{2}^{meson}+Cov_{bb}^{meson})}$ (25)
+$Cov_{bb}^{vol}=\frac{2Cov(p,m)Cov_{bb}(p,m)C_{2}^{meson}-Cov_{bb}^{meson}\left(Cov(p,m)^{2}+Cov_{bb}(p,m)^{2}\right)}{(C_{2}^{meson}-Cov_{bb}^{meson})(C_{2}^{meson}+Cov_{bb}^{meson})}$ (27)
 and
-$C_{2}^{vol,4D}=\frac{-2Cov_{bb}^{meson}Cov_{bb}(p,m)Cov_{bb}(p,m)+C_{2}^{meson}\left(Cov(p,m)^{2}+Cov_{bb}(p,m)^{2}\right)}{(C_{2}^{meson}-Cov_{bb}^{meson})(C_{2}^{meson}+Cov_{bb}^{meson})}\ .$ (26)
-I have verified that the expression (26) for the volume fluctuations contribution to the proton factorial cumulant yields results numerically very close to the simpler expression in Eq.
+$C_{2}^{vol,4D}=\frac{-2Cov_{bb}^{meson}Cov_{bb}(p,m)Cov(p,m)+C_{2}^{meson}\left(Cov(p,m)^{2}+Cov_{bb}(p,m)^{2}\right)}{(C_{2}^{meson}-Cov_{bb}^{meson})(C_{2}^{meson}+Cov_{bb}^{meson})}\ .$ (28)
+I have verified that the expression (28) for the volume fluctuation contribution to the proton factorial cumulant yields results numerically very close to the simpler expression in Eq.
 (16).
@@
 The dashed line represents the estimate of the contribution of volume fluctuations (Eq.
-25).
+27).
 The dotted line represents corresponding correlation coefficient for meson.
-The covariance of proton multiplicities observed at opposite azimuthal angles is nearly independent on the strength of local fluctuations in the deposition of baryons, for $\Delta\phi\leq\pi$ (compare solid and dash-dotted lines in Fig.
+The covariance of proton multiplicities observed at opposite azimuthal angles is nearly independent of the strength of local fluctuations in the deposition of baryons, for $\Delta\phi\leq\pi$ (compare solid and dash-dotted lines in Fig.
 6).
@@
 As the strength of baryon density fluctuations increases, the second factorial cumulant of the proton number increases beyond the volume fluctuation contribution (Fig.
-4), while the covariance of proton multiplicities in two azimuthal angle intervals remains approximately constant (Fig.
+4), while the covariance of proton multiplicities in two opposite azimuthal angle intervals remains approximately constant (Fig.
 7).
@@
 It is preferable to measure separately both the covariance and the second factorial cumulant for protons, along with the associated volume fluctuation contributions and not just the dynamical correlation coefficient (Eq.
-27).
+29).
 An excess of the proton second factorial cumulant over the volume fluctuation baseline, combined with a proton covariance consistent with volume fluctuations, provides strong evidence for short-range fluctuations in the baryon distribution in the transverse plane.
-V Summary and Outlook
+V Conclusions
 The paper presents a method to search for local fluctuations in the fireball produced in relativistic heavy-ion collisions.
 In particular, I study the fluctuations of a specific type of particles of interest (protons) with respect to reference particles (charged mesons).
-The method allows to partly separate global fluctuations of the particle emission from fluctuations on smaller scales.
+The method allows one to partly separate global fluctuations of the particle emission from fluctuations on smaller scales.
 In order to be sensitive to local fluctuations in the transverse plane, I propose to consider fluctuations of particle emission in restricted windows of azimuthal angle.
 Due to the presence of a strong transverse flow in heavy-ion collisions, particles emitted at a given azimuthal angle originate predominantly from a region of the fireball located at that same angle.
-Strong global fluctuations in the particle emission occur due to fluctuations of the collective flow event by event, e.g.
-fluctuations of the magnitude and orientation of the harmonic flow vectors.
-Such global volume fluctuation contributions can be estimated calculating fluctuations of the meson distribution and scaling accordingly.
+Strong global fluctuations in particle emission occur due to fluctuations of the collective flow event by event, e.g., fluctuations of the magnitude and orientation of the harmonic flow vectors.
+Such global volume fluctuation contributions can be estimated calculating fluctuations of meson distributions and scaling accordingly.
 Complementary information is obtained from the covariance of the number of protons emitted in two opposite intervals of azimuthal angle.
@@
 Such a procedure can identify the presence of short range fluctuations in the baryon densities that may carry information about the underlying microscopic dynamics.
-In particular, the method discussed here could be used in the search for fluctuations in baryon density in the vicinity of the QCD critical point [28, 19].
+In particular, the method discussed here could be used in the search for fluctuations of the baryon density in the vicinity of the QCD critical point [28, 19].
 In order to validate the method based on the study of proton distributions in restricted intervals of azimuthal angle further studies are planned.
-These involve : the calculation for net baryons and net protons, using a model with nontrivial higher order proton cumulants, including effects of global baryon conservation, including effects of baryon diffusion, and making predictions for different collisions energies using realistic, 3+1-dimensional hydrodynamic simulations.
-Finally, the dependence of the baryon factorial cumulants on the width of the azimuthal angle window could be potentially sensitive to the value of the correlations length of the underlying baryon fluctuations.
+These involve the calculation for net baryons and net protons, using a model with nontrivial higher order proton cumulants, including effects of global baryon conservation, including effects of baryon diffusion, and making predictions for different collisions energies using realistic, 3+1-dimensional hydrodynamic simulations.
+Using a full 3+1-dimensional dynamics with baryon diffusion could moderate the final effect of baryon density fluctuations.
+Finally, the dependence of the baryon factorial cumulants on the width of the azimuthal angle window could be potentially sensitive to the value of the correlation length of the underlying baryon fluctuations.
 Acknowledgments
-The author thanks Adam Bzdak for useful discussions. This research was supported by the AGH University of Science and Technology and by the Polish National Science Centre grant: 2023/51/B/ST2/01625.
+The author thanks Adam Bzdak for useful discussions.
+This research was supported by the AGH University of Science and Technology and by the Polish National Science Centre Grant No.
+2023/51/B/ST2/01625.
+Data availability
+This manuscript has associated data openly available at [51].